the wave function that describes a certain transverse wave at initial time t=0 is y(x,y)= (7.20mm)sin2π[x•0.3m-t•0.0400 s] 


1
Expert's answer
2021-02-01T15:07:36-0500

a) amplitude


"A=7.2\ mm"

b) wavelength


"\u03bb=0.3m"

c) frequency


"f=\frac{1}{0.04}=25\ Hz"

d) speed of propagation


"v=(25)(0.3)=7.5\frac{m}{s}"

e) period


"T=0.04\ s"

f) wave number


"k=\frac{2\pi}{0.3}=21\ m^{-1}"

g) direction of propagation of the wave is negative "x" direction